OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) -12? x -5 x -5 -5 x A number line with a closed circle on negative 5 and shading to the right. Which of the following is not a way to represent the solution of the inequality 2(x - 1) -12? x -5 x -5 -5 x A number line with a closed circle on negative 5 and shading to the right.

OpenStudy (anonymous):

sorry it double..

OpenStudy (anonymous):

there's no inequality

OpenStudy (anonymous):

idk thats the way the question is

OpenStudy (anonymous):

you forgot to put the < or > signs

OpenStudy (anonymous):

\[\ge \] in front of -12

OpenStudy (anonymous):

2(x-1)>=-12 2x-2>=-12 2x>=-10 x>=-5 closed circle on -5, shade to right

OpenStudy (anonymous):

Thank you